# A Random Rule Model

### 04 Mar 2026 v1

@@ Abstract
We propose a Random Rule Model (RRM) in which behavior is generated by switching among a small library of transparent, parameter-free decision rules. A differentiable gate learns environment-dependent rule propensities, producing an interpretable mixture over named procedures. We develop a global identification theory based on two verifiable conditions on the observed support. Applied to 10,00010{,}000 binary lottery problems, rule-gating substantially outperforms structured neural benchmarks based on expected utility and prospect theory, approaching the most flexible benchmark while remaining highly restrictive under permutation-fit tests, and retains predictive content on an independent dataset. Mechanism diagnostics reveal that extreme-outcome screening, salience, and attention rules carry the largest responsibility weights, with systematic shifts along tradeoff complexity and dispersion asymmetry. Robustness checks confirm that the findings are not driven by the ex-ante library choice, marginal dominance relationships, or the availability of additional regressors.
@@ 1 Introduction
A central task in economics is to represent how behavior varies across environments. In many modern empirical settings, the environment space is vast: researchers observe thousands of distinct decision problems, often with repeated observations for each problem, and the main object of interest becomes the systematic mapping from a decision environment to behavior. One fundamental challenge is therefore not only to fit this mapping accurately - modern machine-learning predictors typically do - but to do so with a representation that remains economically interpretable and empirically disciplined.
This paper proposes a procedural representation of this mapping. Instead of assuming that choices are generated by a single valuation model, we model behavior as switching among a small library of transparent, deterministic decision rules. We call this representation a Random Rule Model (RRM): behavior in a given environment is as if one rule from the library is selected at random and its recommendation is followed. A differentiable gating mechanism assigns propensities to rules as a function of observable features of the environment, so that predicted behavior is an explicitly interpretable mixture over named lines of conduct. In this way, machine-learning optimization is used as a tool to estimate a disciplined economic structure - one that scales to large environments, does not presuppose rationality, and preserves a decomposition of fit into distinct behavioral mechanisms.
A Random Rule Model is attractive as a general representation of choice for at least three reasons. First, it provides an interpretable language for systematic heterogeneity across environments. When behavior changes with the decision problem, the model does not simply register a shift in a latent value function; it states which line of conduct becomes more prevalent and which becomes less so. This makes it possible to summarize large environment-to-behavior mappings in terms of a small set of well-studied behavioral mechanisms - e.g., screening on extremes, focusing on modal outcomes, salience-based comparisons, regret-type comparisons, limited consideration, and related procedures - without committing to a single parametric valuation story that must govern all environments.
Second, RRMs impose a distinctive form of empirical discipline. The model can only fit the mapping from environments to behavior through the patterns of agreement and disagreement generated by the rule library, together with how these patterns vary with observable features. This restriction is neither the knife-edge discipline of a fully specified valuation model nor the unconstrained flexibility of a black-box predictor. It is instead a procedural discipline: the representation is flexible in where different procedures apply, but disciplined in what procedures can be invoked and in how they can combine. This makes RRMs well suited to settings where the environment space is large and rich, but where interpretability requires that predictive success be attributable to a limited repertoire of transparent procedures.
A further advantage is that RRMs do not presuppose the kind of cyclical consistency condition that underlies valuation-based representations. In large menu datasets, such restrictions are often violated, and it is empirically difficult to determine whether violations reflect measurement noise, instability of preferences, or heterogeneity in the latent valuation objects that could rationalize the data. By contrast, the RRM disciplines behavior procedurally. It allows rich non-transitive and context-dependent patterns in the induced choice relation, while keeping interpretation anchored in a transparent library of deterministic rules whose prevalence can vary systematically with observable features of the environment.
This procedural representation raises a foundational question: when is the gate - and hence the behavioral decomposition - pinned down by the data, rather than being a convenient reparameterization of a flexible predictor? We answer this by developing an identification theory for Random Rule Models. The key identifying force is rule switching: because a given rule recommends different sides in different decision problems, it links its latent index across many competitive configurations. We formalize this logic and show how identification reduces to a verifiable property of the realized design: global identification follows from two transparent conditions on the observed menu support: within-feature variation in decisive-side patterns (which rules recommend which side must vary enough at each feature value) and affine richness of the feature values at which this variation occurs. Both conditions are directly verifiable on the data. In our application, the gate operates on a $12$-dimensional vector of interpretable summary statistics $z(A)$ - coarse enough that many distinct menus share the same feature value, making the within-feature conditions empirically testable, while capturing the decision-relevant variation without sacrificing predictive accuracy.
Identification matters because the empirical object of interest is not only predictive accuracy, but an interpretable procedural decomposition: which named procedures are operative, and how their prevalence shifts with primitive features of the environment. The RRM structure makes this decomposition explicit. For each environment, the model first determines which rules are active—i.e., which procedures deliver a strict ranking under the model’s dominance - based decisiveness criterion — then the gate outputs weights across the active rules. Predicted behavior is therefore an explicit mixture of rule-generated recommendations with environment-dependent mixture weights. This structure turns fit into mechanism-level objects: (i) how many rules are typically decisive, (ii) how responsibility is distributed across economically meaningful families of procedures, and (iii) how responsibility shifts with salient features of the environment. Without identification, these decompositions would be fragile; with it, they become disciplined, data-supported statements about procedural variation.
Although the framework applies to general discrete choice environments with arbitrary menu sizes and dominance criteria, we develop this fully in Appendix A, where we state and prove the global identification theorem for menus with finitely many alternatives (Theorem 2). The main text presents the model, identification theory, and diagnostics in the binary choice setting. This is without loss of generality for the conceptual development: the general and binary treatments share the same structure - a grouped-softmax representation, activity-conditional normalization, and verifiable conditions for global identification - with the main difference being notational (log-odds become log-ratios, two-sided menus become multi-sided menus). The binary exposition conveys the essential ideas with lighter notation, and it is also the setting that our empirical application requires.
We bring the framework to a large-scale application in risky choice. We study the choices13k dataset of Peterson et al. (2021), which contains thousands of distinct binary lottery menus with repeated observations per menu, so that each menu can be summarized by a stable choice frequency. This setting is ideal for our purpose: the environment space is rich, predictive accuracy is a meaningful benchmark, and the structure of the mapping from menus to behavior is precisely the object of interest.
We construct a small library of parameter-free decision rules spanning canonical behavioral mechanisms studied in the literature - extreme-outcome screening, modal-outcome comparisons, salience-based comparisons, regret- and disappointment-type comparisons, and limited-consideration-style defaults - and impose an activity discipline under which a rule can influence predicted behavior only when it delivers a decisive recommendation. We estimate a differentiable gate that maps observable menu features into propensities over rules, and we evaluate all models out of sample using the same cross-validation protocol as Peterson et al. (2021). We then use the fitted mixture representation to report rule-level diagnostics (responsibility weights, concentration, and refit-based ablations) and to study how inferred rule responsibility shifts systematically with interpretable menu primitives such as tradeoff complexity and dispersion asymmetry. Finally, we compute and report identification diagnostics for the fitted gate, documenting when and how the procedural decomposition is empirically pinned down on the realized support of menus.
We find that rule-gating delivers large predictive gains relative to structured benchmarks while remaining more disciplined than the most flexible architectures. In out-of-sample evaluation on choices13k, rule-gating (MSE $0.0128$) substantially outperforms Neural EU ($0.0222$), Neural CPT ($0.0215$), Neural PT ($0.0206$), and the value-based network ($0.0192$), and outperforms the context-dependent model ($0.0134$), with only MOT ($0.0114$) achieving lower MSE. Using Neural EU as the baseline and MOT as the flexible benchmark, rule-gating closes about $87\%$ of the baseline-to-MOT gap, compared to $\approx 28\%$ for the value-based model and $7\%$$15\%$ for the prospect-theoretic benchmarks. Under permutation-fit restrictiveness, rule-gating and the structured benchmarks are essentially fully restrictive (near $100\%$), whereas MOT is markedly less so ( ${\approx}\,20\%$), indicating that rule-gating’s high completeness does not come from generic capacity. Portability reinforces this: when exported to CPC18 (Erev et al., 2017; Plonsky et al., 2017) without re-tuning, rule-gating and the context-dependent model transfer well (CPC18 MSE $0.0170$ vs. $0.0172$; trial-level Brier $0.197$ vs. $0.198$), far ahead of the structured benchmarks ( ${\approx}\,0.032$), while MOT deteriorates sharply ($0.0926$).
A central advantage of the RRM is that it does not deliver predictions as an opaque function of menu features. For each menu, the model determines which rules are active—which procedures induce a strict FOSD ranking—and distributes weight across active rules, producing an explicit mixture of rule-generated recommendations. This structure lets us translate predictive performance into interpretable objects: average effective responsibility of each procedure, concentration of the fitted mixture, and refit-based ablations that isolate which procedures are pivotal for prediction.
We find that effective responsibility is concentrated on a small set of psychologically standard procedures: extreme-outcome screening, modal-outcome comparisons, and salience-based comparisons carry the largest non-attention shares. Ablations are even more selective: removing either salience-based comparison produces the largest deterioration, followed by extreme-max and security-type screening, whereas modal-outcome, regret-, and disappointment-type procedures contribute little incremental content. Rule responsibility shifts systematically with objective menu primitives: along tradeoff complexity, the mixture reallocates toward attention and extreme-outcome screening and away from modal-outcome comparisons; along dispersion asymmetry, the pattern reverses on key dimensions.
@@
We subject the framework to a battery of robustness checks addressing identification, library composition, the attention channel, the activity discipline, and the choice of gate features. The identification diagnostics confirm that both conditions of the global identification theorem are satisfied on the Choices13k data, with rule rankings by $w_{f}$ and $\phi(f)$ highly stable across training samples. Cross-fitted library selection retains the bulk of accuracy with $k=7$ rules, and the selected rules coincide with those receiving the highest responsibility weights. Removing the two attention rules brings MSE to $0.0201$, still outperforming most structured benchmarks. Varying the activity discipline reveals that over-restricting the activity set is more costly than removing it, so the FOSD-based activation is best understood as an interpretive discipline rather than a predictive one. A placebo test confirms that the rule indicators carry genuine predictive content, rule-gating outperforms standard classifiers on rule features, and the interpretable gate features $z(A)$ match or improve on the raw $40$-dimensional encoding.
@@
Our approach relates to the literatures on stochastic choice and limited attention or consideration. In random utility models (RUM), choice probabilities are generated by heterogeneity in latent utilities (McFadden, 2006; McFadden and Richter, 1991; Kitamura and Stoye, 2018). A large related literature builds on this framework to account for limited, random, or menu-dependent consideration, jointly recovering preferences and attention/consideration mechanisms from stochastic choice data (Manzini and Mariotti, 2014; Cattaneo et al., 2020; Aguiar et al., 2023). The Random Rule Model (RRM) adopts a different primitive: stochasticity operates through randomization over a small library of interpretable decision procedures rather than through random utilities or random consideration sets. Moreover, the set of procedures that can influence a given menu is disciplined menu-by-menu by an explicit dominance-based activity criterion. The RRM should therefore be viewed as a procedural stochastic-choice model: it provides a complementary, mechanism-oriented description of the choice process that is generally neither nested nor nests, the RUM and attention models.
On the behavioral side, our work complements the literature that explains choice anomalies through menu-dependent perception and evaluation (probability weighting, reference dependence, salience, regret and disappointment, and limited attention; see, e.g., Kahneman and Tversky, 1979; Tversky and Kahneman, 1992; Prelec, 1998; Kőszegi and Rabin, 2006; Bordalo et al., 2012; Loomes and Sugden, 1986; Bell, 1985; Gul, 1991; Masatlioglu et al., 2012; Manzini and Mariotti, 2014). It also connects to the broader tradition of modeling choice as the outcome of procedures and rules under bounded rationality and simplicity constraints (Simon, 1955; Tversky and Kahneman, 1974; Kahneman et al., 1982; Rubinstein, 1988; Aumann, 2008, 2019; Manzini and Mariotti, 2007, 2012; Cherepanov et al., 2013; Rubinstein and Salant, 2006; Salant and Rubinstein, 2008, 2011). Relative to this literature, our contribution is methodological: we provide a scalable estimation strategy for menu-dependent random switching among simple procedures, and we show how to map predictive performance into procedure-level implications in a large dataset.
Our approach builds on the decision-rule framework in Seror (2026). That paper develops a library of decision rules for risky decisions, and introduces concentration and overlap diagnostics to study which rules can rationalize observed choices, without targeting predictive performance. The present paper repurposes the same core primitives to build a predictive Random Rule Model (RRM): behavior is generated by random switching among transparent decision procedures, conditional on which procedures are active on a given menu.
@@
Our mechanism results also relate to the literature on complexity and rule selection (Arrieta and Nielsen, 2024; Halevy and Mayraz, 2024; Puri, 2025). Closest to our analysis, Enke and Shubatt (2023) construct menu-based complexity indices and show that “tradeoff complexity” predicts behavioral signatures in large-scale choice data. We build on this by studying how complexity and dispersion asymmetry are associated with systematic reallocation of procedure weights inside an interpretable random-rule mixture.
@@
Our approach also speaks to the recent agenda of empirical machine learning in economics (Liang, 2025). Our completeness normalization follows recent work benchmarking economic structure against flexible ML baselines (e.g., Peysakhovich and Naecker, 2017; Fudenberg and Liang, 2020; Fudenberg and Puri, 2022; Fudenberg et al., 2022); our restrictiveness diagnostic applies the logic of Fudenberg et al. (2026); and we implement a portability exercise targeting external validity. Finally, a growing literature documents strong out-of-sample performance in risky choice (e.g., Erev et al., 2010, 2017; Plonsky et al., 2016; Peterson et al., 2021). Our contribution is to show that a disciplined mixture of interpretable procedures can recover a large share of these predictive gains while producing mechanism-level objects that are directly interpretable in economic terms.

### 19 Mar 2026 v2

@@ Abstract
We study stochastic choice when behavior is generated by switching among a small library of transparent deterministic decision procedures. The object of interest is procedural heterogeneity: how the relative importance of these procedures varies across decision environments. We model this heterogeneity with a Random Rule Model (RRM), in which menu-level choice probabilities arise from environment-dependent weights on named rules. We show that identification has a two-step structure. At a fixed feature value, variation in decisive-side patterns across menus identifies the vector of relative rule weights up to scale; across sufficiently rich feature values, these recovered weights identify the parameters of an affine gate. Applied to a large dataset of binary lottery choices, the estimated procedure weights are concentrated on a small subset of interpretable rules and shift systematically with menu characteristics such as tradeoff complexity and dispersion asymmetry. Out-of-sample prediction and cross-dataset portability provide supporting evidence that the recovered procedural representation is empirically disciplined.
@@ 1 Introduction
Stochastic choice data are often interpreted through latent utility heterogeneity, random utility shocks, or menu-dependent consideration. An alternative possibility is that observed choice frequencies reflect systematic switching across decision procedures. In many environments, agents may sometimes screen on extreme outcomes, sometimes compare modal consequences, sometimes focus on salient payoff differences, and sometimes follow simpler attention-based procedures. The central object in such settings is not only the average choice probability attached to a menu, but the procedural heterogeneity that determines which decision procedures are more prevalent in which environments.
This paper studies stochastic choice through that lens. We develop a Random Rule Model (RRM) in which behavior is generated by random switching among a small library of transparent, deterministic decision rules. Each rule maps a menu into a recommendation; stochastic choice arises because the relative weight placed on these rules varies with observable features of the menu. The resulting model yields an explicit procedural decomposition of behavior: changes in choice frequencies across environments are interpreted as changes in the prevalence of named lines of conduct rather than as shifts in an unobserved reduced-form predictor.
This representation is attractive for several reasons. First, it provides a tractable language for procedural heterogeneity in large choice environments. Rather than forcing a single valuation model to govern all menus, the model allows different procedures to matter in different regions of the environment space. Second, it remains economically interpretable: the primitive objects are transparent rules with clear behavioral content. Third, it imposes discipline. The model can fit choice frequencies only through the patterns of agreement and disagreement generated by the rule library and through systematic variation in the corresponding rule weights. Fourth, it does not require transitivity or other consistency conditions on the aggregate choice function: because each rule is applied independently and the mixture weights vary freely with menu features, the model can accommodate systematic violations of classical rationality axioms that are well-documented in stochastic choice data.
The rule weights depend on observable features of the menu, and not all rules deliver a strict recommendation on every menu - the model restricts behavioral responsibility to rules that are actually decisive, ensuring that the decomposition attributes weight only where a rule provides a clear-cut ranking. The formal mechanism is a softmax gate over rule-specific indices.
The main theoretical question is whether the latent procedural weights are genuinely identified from menu-level choice frequencies. We show that they are, under a two-layer structure. In the first layer, consider menus that share the same value of the feature vector used by the gate (a possibly coarsened summary of the full menu description). These menus also share the same latent rule weights, but they may differ in which rules recommend which option — some rules switch sides across menus even when the gate features are held fixed. This rule-switching variation pins down the relative importance of the rules at a given set of characteristics: if enough rules switch sides across enough menus, the data uniquely determine how much weight each rule carries. In the second layer, once relative rule weights have been recovered at sufficiently many distinct menu characteristics, the law governing how those weights change with menu features is itself identified. Global identification thus separates cleanly into a within-characteristic step that uses rule switching to recover relative weights, and an across-characteristic step that maps those weights back into the gate.
The identification theorem is constructive: it implies a two-step estimator in which normalized rule weights are first recovered cell by cell from the linear restrictions generated by rule switching, and the gate parameters are then estimated from the cross-cell variation in recovered log-weights. We establish consistency and asymptotic normality for this estimator, yielding inference for economically meaningful functionals of the decomposition.
For the main empirical analysis, however, we estimate the model by minimizing out-of-sample mean squared error (MSE). This avoids the binning of the feature space into discrete cells that the two-step procedure requires and places the rule-gating model on the same predictive footing as the neural-network benchmarks against which it is evaluated. The cellwise identification theory guarantees that the rule weights recovered at each menu characteristic have structural content — they are pinned down by the data, not merely by the predictive objective. When the gate specification is correct, both estimators target the same population decomposition; the two-step estimator provides a built-in diagnostic for this restriction.
We apply the framework to the choices13k dataset of Peterson et al. (2021), which contains thousands of distinct binary lottery menus with repeated observations per menu. We construct a library of twelve parameter-free decision rules spanning canonical behavioral mechanisms — extreme-outcome screening, modal-outcome comparisons, salience-based comparisons, regret- and disappointment-type comparisons, and limited-consideration-style defaults — and estimate the model using the same cross-validation protocol as Peterson et al. (2021). We first verify the identification conditions on the data, then report the MSE-fitted behavioral decomposition: responsibility weights (the average share of predicted choice attributed to each rule), ablation indices (the predictive deterioration when a rule is removed), and concentration measures (how dispersed rule usage is across the library). The two-step econometric estimates follow as a cross-check. We then study how rule responsibility shifts with interpretable menu characteristics such as tradeoff complexity and dispersion asymmetry, and conclude with predictive validation and portability to an independent dataset.
The main empirical findings are as follows. First, on identification: both conditions of the global identification theorem are satisfied on the Choices13k data, with rule rankings by responsibility weight highly stable across training samples and ablation rankings moderately stable. Second, on the behavioral decomposition: the attention channel (left/right defaults) absorbs roughly half of effective responsibility, consistent with substantial inattention in online data collection. Among the structured rules, salience-based comparisons, regret-type comparisons, and modal-outcome screening carry the largest non-attention shares. Ablations are even more selective: removing either salience-based comparison produces the largest deterioration in MSE, followed by regret-type comparisons, extreme-max screening, and modal-outcome comparisons, whereas disappointment-type procedures and the modal-average rule contribute little or negative incremental content. Rule responsibility shifts systematically with objective menu primitives. Along tradeoff complexity, regret shows the steepest decline, the mixture reallocates toward attention and salience-based comparisons, and modal-outcome comparisons weaken, consistent with the view that decision-makers facing more complex tradeoffs fall back on simpler heuristics or default behavior. Along dispersion asymmetry, regret declines steeply while attention defaults rise, suggesting that large differences in spread push the gate toward simpler default behavior.
Third, on predictive performance: in out-of-sample evaluation, rule-gating (MSE $0.0117$) substantially outperforms neural-network implementations of expected utility ($0.0222$), prospect theory ($0.0206$), and cumulative prospect theory ($0.0215$), as well as a value-based network ($0.0192$), and outperforms a flexible context-dependent neural predictor ($0.0134$). Only the Mixture of Theories (MOT, $0.0114$), a mixture-of-experts architecture from Peterson et al. (2021), achieves lower MSE. Rule-gating closes about $97\%$ of the Neural-EU-to-MOT gap. Moreover, rule-gating outperforms MOT at small-to-moderate training fractions; MOT’s advantage emerges only with large samples, consistent with its larger number of free parameters requiring more data to pay off.
Fourth, on restrictiveness and portability: a permutation-fit diagnostic reveals that rule-gating is essentially fully restrictive — when trained on randomly permuted choice frequencies, it does no better than a constant predictor, meaning the model cannot absorb structureless variation. MOT, by contrast, fits permuted targets about five times better than a constant predictor, indicating substantial capacity to absorb structureless variation. This distinction is confirmed by out-of-sample portability: we export all models — without re-tuning — to the Choice Prediction Competition 2018 (CPC18) dataset (Erev et al., 2017; Plonsky et al., 2017), an independent collection of binary lottery choices with different subjects and a different menu sample. Rule-gating and the context-dependent model transfer well, far ahead of the structured benchmarks, while MOT deteriorates sharply.
@@
We subject the framework to a battery of robustness checks addressing library composition, the attention channel, the activity discipline, and the choice of gate features. Cross-fitted library selection retains the bulk of accuracy with $7$ rules, and the selected rules coincide with those receiving the highest responsibility weights. Removing the two attention rules brings MSE to $0.0201$, still close to or outperforming most structured benchmarks. Varying the activity discipline reveals that over-restricting the activity set is more costly than removing it, so the dominance-based activation is best understood as an interpretive discipline rather than a predictive one. A placebo test, in which rule indicators are randomly reassigned across menus while preserving their marginal statistics, confirms that the rule indicators carry genuine, menu-specific predictive content: the gate’s performance depends on the structured correspondence between menus and rule recommendations, not merely on the availability of additional binary regressors. Finally, an interpretable twelve-dimensional vector of summary statistics performs as well as a higher-dimensional raw menu encoding as gate input, indicating that the economically motivated features capture the decision-relevant variation without overfitting.
Related literature.
@@
Our approach relates to the literature on stochastic choice. In random utility models (RUM), choice probabilities are generated by heterogeneity in latent utilities (McFadden, 2006; McFadden and Richter, 1991; Kitamura and Stoye, 2018). A related literature accounts for limited, random, or menu-dependent consideration, jointly recovering preferences and attention/consideration mechanisms from stochastic choice data (Manzini and Mariotti, 2014; Cattaneo et al., 2020; Aguiar et al., 2023). The RRM adopts a different primitive: stochasticity operates through randomization over a small library of interpretable decision procedures rather than through random utilities or random consideration sets. It should be viewed as a procedural stochastic-choice model that is generally neither nested in, nor nests, the RUM and attention frameworks.
Our identification argument connects to the econometrics of latent heterogeneity. The first-stage identification — recovering mixture weights from linear restrictions generated by rule switching — is structurally related to identification in finite mixture models with instrument-like variation (Henry et al., 2014). The novelty is that the identifying variation comes from the discrete patterns of decisive-side recommendations, which are observable and deterministic, rather than from excluded instruments or panel structure.
@@
The paper also connects to the broader tradition of modeling choice as the outcome of procedures under bounded rationality and simplicity constraints (Simon, 1955; Rubinstein, 1988; Aumann, 2008; Manzini and Mariotti, 2007, 2012; Cherepanov et al., 2013; Salant and Rubinstein, 2008), and to the behavioral literature on menu-dependent perception and evaluation (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992; Bordalo et al., 2012; Loomes and Sugden, 1986; Masatlioglu et al., 2012). Our mechanism results relate to the literature on complexity and rule selection (Arrieta and Nielsen, 2024; Halevy and Mayraz, 2024; Puri, 2025; Enke and Shubatt, 2023); closest to our analysis, Enke and Shubatt (2023) construct menu-based complexity indices and show that “tradeoff complexity” predicts behavioral signatures in large-scale choice data. We build on this by studying how complexity and dispersion asymmetry are associated with systematic reallocation of procedure weights inside an interpretable random-rule mixture. Our approach builds on the decision-rule primitives of Seror (2026); the present paper adds the gating mechanism, the identification theory, and the econometric framework.
@@
Finally, our approach also speaks to the recent agenda of empirical machine learning in economics (Liang, 2025). Our completeness normalization follows recent work benchmarking economic structure against flexible ML baselines (e.g., Peysakhovich and Naecker, 2017; Fudenberg and Liang, 2020; Fudenberg and Puri, 2022; Fudenberg et al., 2022); our restrictiveness diagnostic applies the logic of Fudenberg et al. (2026); and we implement a portability exercise targeting external validity. Finally, a growing literature documents strong out-of-sample performance in risky choice (e.g., Erev et al., 2010, 2017; Plonsky et al., 2016; Peterson et al., 2021). Our contribution is to show that a disciplined mixture of interpretable procedures can recover a large share of these predictive gains while producing mechanism-level objects that are directly interpretable in economic terms.